OpenStudy (anonymous):

What is the surface area of a right prism with a HEIGHT of 4, WIDTH of 4, and BASE is 8? 128 in^2 or 160 in^2? If it is 160 in^2, please explain to me how. Please use the SA = 2b + Ph (b is the area of the base P is the perimeter of the base) formula.

OpenStudy (dumbcow):

ok ill use the rectangle w/ dimensions 4 and 8 as my base, so height = 4 Area of Base = 8*4 = 32 Perimeter of Base = 8+8+4+4 = 24 SA = 2*32 + 24*4 SA = 160

OpenStudy (anonymous):

Ah, alright. Thank you :D

OpenStudy (dumbcow):

your welcome